OpenStudy (anonymous):

what is the midpoint of the segment whose endpoints are (0,4) and (-4,-6)?

OpenStudy (black_ninja123):

It will be (-3,0) \[\left( \frac{ 0+(-6) }{ 2 } \right)\frac{ 4+(-4) }{ 2 }\]

OpenStudy (black_ninja123):

@HelpMonster